OpenStudy (anonymous):

standard form of (4+3i)²

OpenStudy (anonymous):

(4+3i)² = (4)² + (3i)² + 2(4)(3i) = 16 + 9i² + 24i Since i² = -1, it becomes = 16 - 9 + 24i = 7+24i

OpenStudy (anonymous):

yes 7+24i